A perfect heat engine takes in heat at 237ºC and rejects heat at 97ºC. What is the efficiency of the engine? How much heat is rejected for every 1000 joules of work done?

1
Expert's answer
2022-04-26T14:56:19-0400

"\eta=1-\frac{97+273}{237+273}=27 .5\%,"

"Q=\frac A\eta=3640~J."